OpenStudy (anonymous):

Triangle ABC with endpoints A(-1,4), B(3,1) and C(0,-3) is a right Triangle. Using the principles of Slope, please verify this.

OpenStudy (anonymous):

Find the gradient of the 3 lines, mAB=(y2-y1)/(x2-x1) =(1-4)/(3--1) =-3/4 mAC=(-3-4)/(0--1) =-7 mBC=(1--3)/(3-0) =4/3 therefore: if mAB*mBC=-1, then AB is perpendicular to BC ,which mean that the angle between AB and BC is 90 degrees..... -3/4*4/3=-1, therefore triangle ABC is right angled